OpenStudy (studybuddie):

299 is what % of 4,600?

Mehek (mehek14):

\(\dfrac{299}{4600}=\) You can use a calculator to find the decimal

undeadknight26 (undeadknight26):

299/4600 =?

undeadknight26 (undeadknight26):

Calculators are very handy devices.

Mehek (mehek14):

After that, move the decimal 2 places to the right

Mehek (mehek14):

That's how you get the percentage

Nnesha (nnesha):

is means equal and the word of means multiply 299= p times 4600 percent out of 100 so\[\rm 299=\frac{P}{100} \cdot 4600\] solve for P